OpenStudy (anonymous):

Analytic Geometry PLS HELP!!! Find an equation for the set of all points whose distance from the origin is twice its distance from x+y-1=0

OpenStudy (anonymous):

Let (u,v) be such a point then \[ \sqrt{u^2+v^2} = 2\frac {|u+v-1|} {\sqrt{1^2 +1^2}}=2 \frac {|u+v-1|} {\sqrt{2}} \]

OpenStudy (anonymous):

Can you finish it now?

OpenStudy (anonymous):

Square both sides and get \[ u^2 + v^2 = 2 (u+v-1)^2 \]

OpenStudy (anonymous):

\[ u^2+4 u v-4 u+v^2-4 v+2=0 \]

OpenStudy (anonymous):

if I use variables x and y: is x^2=y^2-4x-4y-4=0 the answer?

OpenStudy (anonymous):

should be x^2+y^2-4x-4y-4=0

OpenStudy (anonymous):

\[ 2 - 4 x + x^2 - 4 y + 4 x y + y^2=0 \]

OpenStudy (anonymous):

It is a hyperbola. I will post the graph soon.

OpenStudy (anonymous):

I just can't understand the thing about the origin... cause we're given a formula... and it goes like this: (Ax+By+C) / +or - sqrt (A^2+B^2)

OpenStudy (anonymous):

Here you can see the line and the graph of the hyperbola. You can see the distance from a point on the hyperbola is twice the distance to the line.

OpenStudy (anonymous):

What is the distance from (u,v} to the origin (0,0)?

OpenStudy (anonymous):

\[ (u-0)^2 + (v-0)^2=u^2 + v^2 \] This is the distance formula between two points.

OpenStudy (anonymous):

The other is the distance from (u,v) to a line a x + b y + c =0 \[ d=\frac{| au + bv + c|} {\sqrt{a^2 + b^2}} \]

OpenStudy (anonymous):

I tried to solve it and I got x^2+y^@+4xy-4y-4x+2=0 is it the right answer?

OpenStudy (anonymous):

y^2*

OpenStudy (anonymous):

yeah and that's what I got after combining similar terms

OpenStudy (anonymous):

This is the right answer \[2 - 4 x + x^2 - 4 y + 4 x y + y^2=0 \]

OpenStudy (anonymous):

It seems that you got it right

OpenStudy (anonymous):

yeah it's the same thing...I just arranged it

OpenStudy (anonymous):

woaAh thanks for your help ^^

OpenStudy (anonymous):

Did you see the nice graph, I drew it for you?

OpenStudy (anonymous):

I just saw it... you need not to draw it... but thanks for the great effort ^^

OpenStudy (anonymous):

Just another confirmation that our work was right